If 9x* and (5x+62)* are the measures of complementary angles, what is the measure of each angle?
Answer by stanbon(48522)

If 9x* and (5x+62)* are the measures of complementary angles, what is the measure of each angle?
---------------
9x + 5x+62 = 90
14x = 28
x = 2
---------
1st angle: 9x = 18
2nd angle: 5x+62 = 10+62 = 72
===============================
Cheers,
Stan H.